# Photo-nuclear cross sections on Au

### v1

@@ IV Experiment
Targets were activated at seven different central photon beam energies (see Table 3). The gamma energy at the outermost ring typically was close to the central energy of the next lower energy irradiation, thus providing an internal consistency check for these measurements. Using targets that of three concentric segments, the cross-section measurements for the reactions Au(,n)Au and Au(,3n)Au were determined for 21 different energies. The target stack was irradiated for 8-12 hours by a photon beam with fluxes that ranged between 10-10 /s, limited by a copper pre-collimator. The boundaries of the photon beam of 30-40 mm in diameter were shaped by the copper apertures inside the FEL aperture systems and by the copper pre-collimator. Irradiation times, beam intensities and uncertainties are summarized in Table 3.
@@ IV.1 Target stack
The target stack was composed of layers of Au, TiO, Zn, Os and Au as shown in Figure 5 and listed in Table 4. The target segments were mounted in 3D-printed acrylic with cavities made for receiving powdered target material. The cavities were filled with a known mass of the target material and then sealed by adding and curing 3D printing gel on top of the target material. The effects of photon attenuation and scattering by the low-acrylic target mounts were negligible. The thickness of the targets varied across the numerous targets made for this work (see Table 4). The front and back Au targets were arranged in this manner in anticipation of using these to normalize the beam flux and calibrate the beam energy throughout the stack. Placed front to back in the stack, relative to the incident beam, and between the front and back Au targets, were TiO, Zn and Os ones. The beam flux and scattered gamma contribution to the cross section were calculated. There were 3 sets of targets to enable reuse after sufficient decay of the activations in subsequent runs separated by months.
@@ IV.2 Photon beam flux
A significant portion of the overall systematic uncertainty in the cross sections measured in this work is due to the errors in measuring the absolute -ray beam flux. A schematic diagram of the system used to measure this flux during target irradiation is given in Figure 6. The beam flux was measured using three detectors during target irradiation: 1) a shielded plastic scintillator paddle that detects rays and recoil electrons from Compton scattering of the beam passing through the mirror of the optical cavity (labeled MP in Figure 6), 2) a thin plastic scintillator centered in the beam path just after the collimator (labeled SP in Figure 6), and 3) a silicon detector that measures charged particles emitted from photon-induced reactions in a thin CD foil inside an evacuated scattering chamber (labeled CP in Figure 6). The MP detector monitors a portion of the full -ray beam intensity passing through the cavity mirror. The detection efficiency of MP for measuring the beam passing through the collimator is determined using Molly (see Figure 6). This detector can be positioned in the beam path, and a system of high-precision copper absorbers can be inserted into the beam to attenuate the incident flux on Molly to keep the data acquisition system deadtime lower than 40 and the pulse pileup rate of the detector signal below 10. The efficiencies of the other flux monitors, SP and CP, are calibrated relative to the MP using the full beam intensity. The efficiency of MP is determined using the following relationships,
@@
where . The symbol is the photon attenuation coefficient (cm/g) averaged over the beam intensity profile, is the density of the copper attenuators (8.96 g/cm), and x is the length in cm of the copper attenuator inserted in the beam path to reduce the flux incident on Molly. The value of was calculated using the photon attenuation coefficients as a function of energy, (E), obtained from NIST [38, 39], and the beam intensity profile simulated as described in [40]. An example of and , the unattenuated and attenuated beam intensity profiles, used to calculate at the peak beam energy, = 18.0 MeV, is plotted in Figure 7. The value is the -ray energy, E, where the distribution has a maximum value. The distribution is normalized such that the integration over the distribution is equal to 1.0. Because the uncertainty in has a significant impact on the error of the computed -ray transmission through the copper absorbers, the calculated value was benchmarked at several beam energies with in-situ measurements. The average variance between the calculated and measured values was 1.0. This variance was assigned as an estimate of the uncertainty in the calculated values. In turn, the uncertainty in was used to compute the average error in the calculated intensity. The main systematic uncertainty in the measurement of the -ray beam flux arises from the error in determining the effective x-ray attenuation coefficient, , used to compute the -ray transmission through the copper absorbers. The systematic uncertainty in the -ray beam flux determination is 11.0.
IV.3 -ray spectra
@@
After activation, -ray spectra were measured using 12 High Purity Germanium (HPGe) detectors, each with a co-axial cylindrical crystal positioned with its front face 5 cm from the activated target. The measurements started approximately 30 minutes after the end of bombardment and lasted long enough to obtain sufficient counts for the second isomeric state of Au with the half-life of 9.6 hours. Each HPGe detector was calibrated with a mixed calibration source, and the detector efficiencies for the gold isotopes were determined by the given formula [42].
@@
where x and y are ln and ln/1000), respectively. E is the -ray energy from selected nuclei. The absolute efficiency at 333 keV for the Au transition and the corresponding uncertainty is 1-3 and 2, depending on the HPGe being considered. During the measurements, the dead time was typically 1-2. The spectra at the photon beam energies of 18 and 27 MeV are presented in Figure 8. The emitted -ray peaks of 147.81, 188.27 keV (second isomeric state of Au), 333.03, 355.73 keV (ground state of Au) and 293.55, 328.47 keV (Au) are visible in Figure 8.
@@ V Cross-section analysis
The photonuclear cross sections were determined using equations 4 and 5. In addition, the detector efficiency, detector dead time, branching ratio, beam attenuation and self-absorption effects are taken into account. With these corrections, the cross sections can be expressed as,
@@
where and are the cross sections for the ground + 1 isomeric state and 2 isomeric state, respectively. stands for the detected counts, and are the beam flux and the detector efficiency, respectively. The total correction factor above is expressed as = . The , , , and symbols are corrections for the detector efficiency, detector dead time during the measurements, -ray branching ratio, -ray beam attenuation and self-absorption by passing through the targets, respectively. Correction of the photon beam flux due to attenuation, and of the emitted -ray yield due to self-absorption are taken into account by,
@@
where and x are attenuation coefficients for the beam and emitted -ray and the total length of the target, respectively. The statistical error was calculated by RADWARE software [42] and three systematic errors for the -ray branching ratio, detector efficiency and beam flux were determined. Therefore, the total experimental error can be obtained by,
@@
where , , and are uncertainties from the number of nuclei measured by counts, photon beam flux, HPGe detector efficiency and the branching ratio of the selected nuclei, respectively.
@@
After activation by the photon beam, the segments in the target stacks were separated and counted individually with HPGe detectors. The perceptible Au radioisotopes produced via Au(,xn) were Au and Au. The Au nucleus was observed to produce two isomeric states (m= 5, m= 12) and the ground state (2). The simplified decay schemes for the Au and Au nuclei are illustrated in Figure 9. The half-life of the , and ground states of Au are 8.1 s, 9.6 h and 6.17 d, respectively. The m2 and ground states are measurable and distinguishable by counting the rays from the gold targets within a few hours from the end of each irradiation period. The m isomeric state, with its short half-life of only 8.1 s, has its yield included in the decay of the ground state. The 147.81 and 188.24 keV -ray lines shown in Figure 9 and the peaks in Figure 8 are used to determine the yield of the m isomer, and the 333.03, 328.47 keV -ray transitions following emission from ground state of Au are used to calculate the cross section for the sum of the ground state and m1 isomer, respectively. The ground state (1) of Au has a 38.02 h half-life and decays by emission to Pt. The 293.55 and 328.47 keV -ray peaks were used to determine the cross section of the Au(,3n)Au reaction. The decay information for the Au and Au nuclei, along with uncertainties, is listed in Table 5.
@@
| nucl. | half-life | peak (keV) | branch ratio () |
| Au | 6.1669 (6) d | 333.03 | 22.9 (9) |
| 355.73 | 87 (3) | ||
| Au | 9.6 (1) h | 147.81 | 43.5 (15) |
| 188.27 | 30.0 (15) | ||
| Au | 38.02 (10) h | 293.549 | 10.9 (3) |
@@
V.1 Au(,n)Au
The cross sections of the combined ground and first isomeric states and second isomeric state for the photon-induced reaction on Au are obtained with the emitted rays of 333.03, 355.73 keV and 147.81, 188.27 keV, respectively. The excitation functions for the Au(,n)Au and Au(,n)Au reactions are presented in Figure 10 together with the previous experimental data [27, 28, 35, 34, 32, 33, 31, 30, 29] and theoretical calculations from TENDL 2021 and JENDL 4.0. Here, the cross sections in Figure 10 were determined by means of weighted mean and error. The cross sections measured in this work with statistical and total uncertainties are given in Table 6.
| nucl. | Au | Au | |||||||||||||||
@@ V.2 Au(,3n)Au
The cross sections for the Au(,3n)Au reaction are obtained from the emitted rays of 293.55 and 328.47 keV. The weighted mean cross sections are presented in Figure 11, together with the previous experimental data [31] and theoretical calculations from TENDL 2021 and JENDL 4.0. The cross sections measured in this work with statistical and total uncertainties are summarized in Table 7.
| nucl. | Au | ||||||||
@@ VI Results and Discussion
This work demonstrates the successful development of a novel and highly efficient method for concurrently measuring photonuclear reaction cross sections at multiple discrete photon energies for various materials using a single irradiation. As a proof-of-concept, the excitation function of the Au(,n)Au and Au(,3n)Au reactions were measured with this novel method utilizing target stacks consisting of concentric ring segments that were bombarded with an uncollimated beam with well-defined radial energy and flux distributions. The measurements were performed at 21 discrete beam energies over the incident photon beam energy range of 13-31 MeV (see Table 6). For the first time, the cross sections of the combined ground and first isomeric states at the beam energy above 20 MeV and the second isomeric state for the Au(,n)Au reactions were determined via the activation method which yielded exclusive cross sections which are unaffected by multiple neutron emission. The method developed here based on radial target segmentation and target stacking makes efficient use of beam time by concurrent irradiation of multiple targets at multiple discrete -ray beam energies. Our Au(,n) cross-section data for the production of Au in the ground state are in agreement with previous data measured using mono-energetic photon beams, but differ from data measured with a Bremsstrahlung beam above 14 MeV (see Figure 10 and Table 6). This observation is striking and deserves further investigation in future work. Furthermore, it can be seen that the evaluations based on TENDL and JENDL significantly underpredict the cross sections for energies above the giant dipole resonance (15MeV). This observation is likely to impact flux calibrations based on Au activation cross sections at these higher energies. In addition, experiments extending the measured cross sections further would provide further guidance to improve evaluations. Our data for the Au(,3n) reaction are given in Table 7 and plotted in Figure 11 in comparison to previous data and model calculations. As can be inferred from Figure 11, additional data are desirable to better evaluate the predictive power of the TENDL and JENDL approaches.
As discussed in Section I, the (,p) reaction is the primary pathway for producing high-specific-activity radioisotopes with electron LINACs. Our cross-section analysis can be extended to the remaining targets (TiO, Zn, and Os) to obtain (,p) reaction cross section. The main reactions of interest for the remaining targets are Ti(,p)Sc, Zn(,p)Cu, and Os(,p)Re. The data obtained for these reactions will be available in future publications.
@@ VI.1 Incident photon scattering effect
The PHITS predictions of the scattering of the incident photons are presented in Figure 12. A few percentage points of the incident high-energy photons from the original beam are lost or scattered due to interactions with the atomic electrons in the target stack (Compton scattering), producing a lower-energy photon tail (see red histogram in Figure 12). Thus, the back Au target has an apparent 3 higher cross section than the first Au target in the beam energy range of 20-27 MeV. The effect increases the apparent cross section in the high-energy region by 2 mb, but has a very small effect in the giant resonance region where the cross sections are approximately 100 times larger, as shown in the simulations of Figure 12(b). Cross sections for the other targets in the stack will be corrected for this effect in future publications.

### v2

@@ IV Experiment
Targets were activated at seven different central photon beam energies (see Table 3). The gamma energy at the outermost ring typically was close to the central energy of the next lower energy irradiation, thus providing an internal consistency check for these measurements. Using targets with three concentric segments, the cross-section measurements for the reactions 197Au(,n)196Au and 197Au(,3n)194Au were determined for 21 different energies. The target stack was irradiated for 8-12 hours by a photon beam with fluxes ranging between 108-109 /s, limited by a copper pre-collimator. The boundaries of the photon beam of 30-40 mm in diameter were shaped by the copper apertures inside the HIS aperture systems and by the copper pre-collimator. Irradiation times, beam intensities and uncertainties are summarized in Table 3.
@@ IV.1 Target stack
The target stack was composed of layers of Au, TiO2, Zn, Os and Au as illustrated in Figure 4 and listed in Table 4. The target segments were mounted on 3D-printed acrylic with cavities made for receiving powdered target material. The cavities were filled with a known mass of the target material and then sealed by adding and curing 3D printing gel on top of this target material. The effects of photon attenuation and scattering by the low-acrylic target mounts were negligible. The thickness of the targets varied across the numerous ones made for this work (see Table 4). The front and back Au targets were arranged in this manner in anticipation of using these to normalize the beam flux and calibrate the beam energy throughout the stack. Placed front to back in the stack, relative to the incident beam, and between the front and back Au targets, were TiO2, Zn and Os ones. The beam flux and scattered gamma contribution to the cross section were calculated. There were 3 sets of targets to enable reuse after sufficient decay of the activations in subsequent runs separated by several months.
@@ IV.2 Photon beam flux
A significant portion of the overall systematic uncertainty in the cross sections measured in this work is due to the errors in measuring the absolute -ray beam flux. A schematic diagram of the system used for this purpose during target irradiation is given in Figure 5. The beam flux was measured using three detectors during target irradiation: 1) a shielded plastic scintillator paddle which detects rays and recoil electrons from Compton scattering of the beam passing through the mirror of the optical cavity (labeled MP in Figure 5), 2) a thin plastic scintillator centered in the beam path just after the collimator (labeled SP in Figure 5), and 3) a silicon detector which measures charged particles emitted from photon-induced reactions in a thin CD2 foil inside an evacuated scattering chamber (labeled CP in Figure 5). The MP detector monitors a portion of the full -ray beam intensity passing through the cavity mirror. The absolute detection efficiency of MP for measuring the beam passing through the collimator is determined using Molly (see Figure 5). The Molly detector is described earlier. Photon interactions in Molly are modeled with GEANT-4. According to these GEANT-4 simulations, the full-energy peak efficiency of Molly is greater than 98 at energies above 10 MeV. Molly can be positioned in the beam path, and a system of high-precision copper absorbers can be inserted into the beam to attenuate the incident flux on Molly to keep the data acquisition system deadtime lower than 40 and the pulse pileup rate of the detector signal below 10. The efficiencies of the other flux monitors, SP and CP, are calibrated relative to MP using the full beam intensity. In this arrangement, the ratios of the rates in SP and CP to that in MP are computed to monitor the stability of the fraction of the total -ray flux that is transmitted through the collimator. Over the duration of each run, the efficiency of MP is determined using the following relationships,
@@
where . The symbol is the photon attenuation coefficient (cm2/g) averaged over the beam intensity profile, is the density of the copper attenuators (8.96 g/cm3), and x is the length in cm of the copper attenuator inserted in the beam path to reduce the flux incident on Molly, while is the net counting rate in the MP paddle. The symbol denotes the -ray energy where the distribution has a maximum value, see Figure 6. The value of was calculated using the photon attenuation coefficients as a function of energy, (E), obtained from NIST [36, 37], and the beam intensity profile simulated as described in [27]. An example of and , the simulated unattenuated and attenuated beam intensity profiles, used to calculate for the beam energy distribution with the peak beam energy, = 18.0 MeV, is plotted in Figure 6. The value is the -ray energy, E, where the distribution has a maximum value. The distribution is normalized such that the integration over the distribution is equal to 1.0. The main systematic uncertainty in determining in the above equation is due to the error in the value of the attenuation . The relative uncertainty in the measured is less than 1 and is not included in the derivation of . The relative uncertainty in is,
The calculated value was benchmarked at several beam energies with in-situ measurements. The average variance between the calculated and measured values was 0.7. We use this variance as an estimate of the uncertainty in the calculated value. In turn, the uncertainty in was used to compute the average error in the calculated intensity. For these measurements, the relative systematic uncertainty in the -ray beam flux is = (0.007)(8.96 g/cm3)(0.033 cm2/g)(39 cm) = 0.081 = 8.1. The relative uncertainty in transfers to the uncertainty in measured MP efficiency, see equation 5. The statistical uncertainty in is less than 1. The efficiency of MP as a function of is displayed in Figure 7. The error bars are the combined statistical and systematic uncertainties.
IV.3 The activated target -ray decay spectra
@@
After activation, -ray spectra were measured with 12 High Purity Germanium (HPGe) detectors, each with a co-axial cylindrical crystal positioned with its front face 5 cm from the activated target. The measurements started approximately 30 minutes after the end of bombardment and lasted long enough to obtain sufficient counts for the detection of the second isomeric state of 196Au with the half-life of 9.6 hours. Typical HPGe detector efficiency shown in Figure 8 was calibrated with a mixed calibration source, and the detector efficiencies for the gold isotopes were determined by the formula [39].
@@
where x and y are ln and ln/1000), respectively, Eγ is the -ray energy from selected nuclei. The absolute detection efficiency at 333 keV for the 196Au transition is 1-3, and the corresponding uncertainty, is 2. During the measurements, the dead time was typically less than 2. The spectra at the photon beam energies of 18 and 27 MeV are presented in Figure 8. The -ray peaks of 147.81, 188.27 keV (second isomeric state of 196Au), 333.03, 355.73 keV (ground state of 196Au) and 293.55, 328.47 keV (194Au) are visible in Figure 8.
@@ V Cross-section analysis
The photonuclear cross sections were determined using equations 14 and 15. In addition, the detector efficiency, detector dead time, branching ratio, beam attenuation and self-absorption effects are taken into account. With these corrections, the cross sections can be expressed as,
@@
where and are the cross sections for the ground + 1st isomeric state and 2nd isomeric state, respectively. stands for the detected counts, and are the beam flux and the detector efficiency, respectively. The total correction factor above is expressed as = . The , , , and symbols are corrections for the detector efficiency, detector dead time during the measurements, -ray branching ratio, -ray beam attenuation and self-absorption by passing through the targets, respectively. Correction of the primary photon beam flux due to attenuation, and of the emitted -ray yield from Au isotopes due to self-absorption are taken into account by,
@@
where and x0 are attenuation coefficients for the beam and emitted -ray and the total length of the target, respectively. The statistical error was calculated by the RADWARE software [39] and three systematic errors for the -ray branching ratio, detector efficiency and beam flux were determined. Therefore, the total experimental error can be obtained by,
@@
where , , and are uncertainties from the number of nuclei measured by counts, photon beam flux, HPGe detector efficiency and the branching ratio of the selected nuclei, respectively. The uncertainties of the branching ratios are given in Table 5.
@@
After activation by the photon beam, the segments in the target stacks were separated and counted individually with HPGe detectors. The perceptible Au radioisotopes produced via 197Au(,xn) were 196Au and 194Au. The 196Au nucleus was observed to produce two isomeric states (m1= 5+, m2= 12-) and the ground state (2-). The simplified decay schemes for the 196Au and 194Au nuclei are illustrated in Figure 9. The half-life of the , and ground states of 196Au are 8.1 s, 9.6 h and 6.17 d, respectively. The m2 and ground states are measurable and distinguishable by counting the rays from the gold targets within a few hours from the end of each irradiation period. The m1 isomeric state, with its short half-life of only 8.1 s, has its yield included in the decay of the ground state. The 147.81- and 188.24-keV -ray lines shown in Figure 9 and the peaks in Figure 8 are used to determine the yield of the m2 isomer, and the 333.03-, 328.47-keV -ray transitions following emission from the ground state of 196Au are used to calculate the cross section for the sum of the ground state and m1 isomer, respectively. The ground state (1-) of 194Au has a 38.02 h half-life and decays by emission to 194Pt. The 293.55 and 328.47 keV -ray peaks were used to determine the cross section of the 197Au(,3n)194Au reaction. The decay information for the 196Au and 194Au nuclei, along with uncertainties, is listed in Table 5.
@@
| nucl. | decay mode | half-life | energy | branching ratio |
| (keV) | () | |||
| 196Aug | 93.0 | 6.1669 (6) d | 333.03 | 22.9 (9) |
| 355.73 | 87 (3) | |||
| 196Aum2 | IT 100 | 9.6 (1) h | 147.81 | 43.5 (15) |
| 188.27 | 30.0 (15) | |||
| 194Aug | 100 | 38.02 (10) h | 293.549 | 10.9 (3) |
@@
V.1 197Au(,n)196Au
The cross sections of the combined ground and first isomeric states and second isomeric state for the photon-induced reaction on 197Au are obtained with the emitted rays of 333.03, 355.73 keV and 147.81, 188.27 keV, respectively. The excitation functions for the 197Au(,n)196Au and 197Au(,n)196Au reactions are presented in Figure 10 together with the previous experimental data [24, 25, 34, 33, 31, 32, 30, 29, 28] and theoretical calculations from TENDL 2021 and JENDL 4.0. Here, the cross sections in Figure 10 were determined by means of weighted mean and error. The cross sections measured in this work with statistical and total uncertainties are given in Table 6.
| nucl. | 196Au | 196Au | |||||||||||||||
@@
The cross sections for the 197Au(,3n)194Au reaction are obtained from the emitted rays of 293.55 and 328.47 keV. The weighted mean cross sections are presented in Figure 11, together with the previous experimental data [30] and theoretical calculations from TENDL 2021 and JENDL 4.0. The cross sections measured in this work with statistical and total uncertainties are summarized in Table 7.
| nucl. | 194Au | ||||||||
@@ VI Results and Discussion
This work demonstrates the successful development of a novel and highly efficient method for concurrently measuring photonuclear reaction cross sections at multiple discrete photon energies for various materials using a single irradiation. As a proof-of-concept, the excitation function of the 196Au(,n)196Au and 196Au(,3n)194Au reactions were measured with this novel method utilizing target stacks consisting of concentric ring segments bombarded with an uncollimated beam with well-defined radial energy and flux distributions. The measurements were performed at 21 discrete beam energies over the incident photon beam energy range of 13-31 MeV (see Table 6). For the first time, the cross sections of the combined ground and first isomeric states at beam energies above 20 MeV and the second isomeric state for the 197Au(,n)196Au reactions, were determined via the activation method which yielded exclusive cross sections which are unaffected by multiple neutron emission. The method developed here based on radial target segmentation and target stacking makes efficient use of beam time by concurrent irradiation of multiple targets at multiple discrete -ray beam energies. Our 197Au(,n) cross-section data for the production of 196Au in the ground state are in agreement with previous data measured using mono-energetic photon beams, but differ from data measured with a bremsstrahlung beam above 14 MeV (see Figure 10 and Table 6). Our data, which were measured with a mono-energetic -ray beam using activation techniques, are in good agreement with data measured with a similar beam but using neutron counting techniques. This result suggests that the standing discrepancies between data are most likely associated with beam type and not experimental techniques, i.e., neutron counting versus target activation. Furthermore, the TENDL and JENDL evaluations significantly underpredict the cross sections at energies above 18 MeV, suggesting a need to update these evaluations. Our data for the 197Au(,3n) reaction are given in Table 7 and plotted in Figure 11 in comparison with previous data and model calculations. As can be inferred from Figure 11, additional data at energies above 30 MeV are desirable to better estimate the predictive powers of the TENDL and JENDL evaluations. As discussed in Section I, the (,p) reaction is the primary pathway for producing high specific-activity radioisotopes using bremsstrahlung rays generated by electron LINACs. Our cross-section analysis can be extended to the remaining targets (TiO2, Zn, and Os) to obtain (,p) reaction cross sections. The main reactions of interest for the remaining targets are 48Ti(,p)47Sc, 68Zn(,p)67Cu, and 189Os(,p)188Re. The data obtained for these reactions will be reported in future publications.
@@ VI.1 Incident photon scattering effect
Having accurate values for the beam flux at each segment of each target is essential in determining the corresponding reaction cross sections. The fractional photon beam fluxes at each target segment were obtained from simulations, which were benchmarked with measurements as described above. With these incident beam fluxes on the segments of the front target, the beam intensity at each target was calculated using the beam transport simulation performed by the PHITS Monte-Carlo code [41]. The geometry illustrated in Figure 4 was implemented in the simulations. The targets in the stack were Au, TiO2, Zn, Os and Au. Successive targets were separated by 0.45 cm. The detailed information about the segments of the actual target stack can be found in Section IV. A total of 15 segments in the target stack (3 segments per target) were exposed to the photon beam at once. The beam flux was simulated for each irradiation (See Table 8). In the energy range of our measurements, Compton scattering and pair production dominate the beam attenuation and scattering as it passes through the stack. The attenuation of the gamma beam is simulated in detail using the PHITS code with the overall attenuation predicted to be 10. The central beam energy, one-sigma radial energy spread, and fractional beam flux on each segment of the stack are given in Table 8. The overall simulation is validated using the data from the gold target at the back of the stack when compared with the data from the one in the front. Predictions of distortions to the photon energy spectrum due to the photons scattering as they pass through the target stack are made using PHITS. Some results from the PHITS calculations are found in Figure 12. In these calculations is about 25 MeV, as can be inferred from the shaded peak in Figure 12 a). As the photon bunch traverses the target stack, the energy of a small fraction of the photons is degraded due to Compton scattering from atomic electrons. This effect creates the low-energy tail on the photon energy distribution (red curve) displayed in Figure 12 a). Because this low-energy tail extends through the GDR region, the cross section measured with the back Au target is higher than that measured using the front one in the beam energy range of 20-27 MeV. This effect increases the apparent cross section in the high-energy region by 2 mb, but has a very small effect in the giant resonance region where the cross sections are approximately 100 times larger, as shown by the simulations plotted in Figure 12 b). Cross sections for the other targets in the stack will be corrected for this effect in future publications.